OpenStudy (anonymous):

Rewrite the function in intercept form. g(x)=3x^2-8x+5

OpenStudy (anonymous):

g(x)=3(x^2-8/3x+16/9) +5+16/3 .....g(x)=3(x-4/3)^2+31/3